What is the mass in grams of 1 mole of Uranium Dioxide (UO2)?

A. 254 g
B. 270 g
C. 540 g
D. 1080 g
E. None of the Above
1
Expert's answer
2018-01-14T07:05:07-0500
m(UO2)=n(UO2)xM(UO2),
M(UO2) - molar mass Uranium Dioxide
M(UO2)= M(U)+ M(O2)=238+16x2=270 g/mole
if n(UO2)=1 mole
then m(UO2)=1x270=270 g
Correct answer – “B”.